# Application of Long-Short Term Memory and Convolutional Neural Networks for Real-Time Bridge Scour Forecast

### v1

@@ Abstract
Scour around bridge piers is a critical challenge for infrastructures around the world. In the absence of analytical models and due to the complexity of the scour process, it is difficult for current empirical methods to achieve accurate predictions. In this paper, we exploit the power of deep learning algorithms to forecast the scour depth variations around bridge piers based on historical sensor monitoring data, including riverbed elevation, flow elevation, and flow velocity.
We investigated the performance of Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN) models for real-time scour forecasting using data collected from bridges in Alaska and Oregon from 2006 to 2021. The LSTM models achieved mean absolute error (MAE) ranging from 0.1m to 0.5m for predicting bed level variations a week in advance, showing a reasonable performance. The Fully Convolutional Network (FCN) variant of CNN outperformed other CNN configurations, showing a comparable performance to LSTMs with significantly lower computational costs. We explored various innovative random-search heuristics for hyperparameter tuning and model optimisation which resulted in reduced computational cost compared to grid-search method. The impact of different combinations of sensor features on scour prediction showed the significance of the historical time series of scour for predicting upcoming events. Overall, this study provides a greater understanding of the potential of Deep Learning (DL) for real-time scour forecasting and early warning in bridges with diverse scour and flow characteristics including riverine and tidal/coastal bridges.
@@
Scour Prediction, Deep-learning, Time Series Forecasting, Hyperparameter Tuning, Long-short Term Memory Network, Convolutional Neural Network
@@ 1 Introduction
Artificial Intelligence (AI) and Machine Learning (ML) for scour depth estimation have been investigated by a range of algorithmic strategies, such as evolutionary computing, fuzzy logic, artificial neural networks, support vector machines, and decision trees (see section  2). These proposed techniques have shown promising results in estimating maximum scour depth, often outperforming traditional empirical equations. However, the performance of these models is still limited by the availability and quality of training data, which are often scarce and may not cover a wide range of geological, hydraulic and geomorphological conditions. Furthermore, these models are typically trained to predict a maximum scour depth for a given flow discharge and cannot be used for dynamic scour prediction and real-time forecast.
Yousefpour et al. phase1 ; phase2 pioneered the application of deep learning (DL) solutions, using Long Short Term Memory (LSTM) networks hochreiter1997long for real-time scour forecast. They leveraged the LSTM’s superiority in temporal pattern recognition and capturing the underlying physics without direct feature extraction. They used historical scour monitoring data in Alaska to train DL models and predict future bed elevation variations and upcoming scour events. Fig. 1 shows the process of local scour around a bridge pier; scour happens when sediments are eroded and washed away by strong vortices in interaction with the bridge substructure, often accelerating during floods. Stage (flow/water level) and sonar (bed elevation) sensors are commonly used sensors installed on the bridge piers to monitor scour by collecting real-time flow and bed elevation data. The variants of LSTM models developed in phase2 showed promising performance in forecasting the scour depth a week in advance, for case study bridges in Alaska. However, there remain research questions about the capability of these algorithms to be upscaled for implementation in practice and deployment in different locations with variable meteorological, geological, geomorphological, and flow conditions.
@@
In this study, we addressed fundamental research questions about the efficiency and generalization of the LSTM scour forecast models in diverse temporal and geographical domains. The LSTM models are compared against temporal CNN as an alternative deep learning solution, in terms of performance and computational cost. Random-Search (RS) heuristics for hyperparameter (HP) tuning are evaluated  RS1 to explore alternate computationally efficient methods to Grid-Search (GS) used in  phase2 for model optimisation. We also investigated the impacts of monitored features (input parameters), including stage, sonar, and velocity (discharge), on the performance of scour forecast models. To summarise, four research questions were addressed in this study:
@@
(RQ2) How do other competitor models, including CNN with vanilla and dilated convolutions, compare with LSTMs for scour forecast?
@@
We carefully designed experiments and developed heuristics to investigate the aforementioned research questions as discussed in the following sections.
@@ 2.2 Real-time Scour Monitoring and Forecast
Although the worst-case scenario (usually a 1-in-100-year flood) is considered for estimating maximum scour depth in bridge design, many uncertainty factors could still lead to scour-related bridge failures: Larger flood events can occur due to extreme weather conditions and climate change; complexities in the scour process, particularly sedimentation and riverbed soil behaviour in the interaction with the flow which can result in larger scour depth; older bridges could fall behind the best recommended practices for scour estimation khandel2019integrated ; lamb2017vulnerability . Considering all these factors, the transport authorities in the past two decades have been resorting to regular monitoring to manage scour, including regular inspections and surveys as well as advanced remote sensing.
In recent years, real-time scour monitoring sensors have become a prominent solution installed on many critical and large-scale bridges lagasse1997instrumentation ; briaud2011realtime ; lebbe2014failure ; PRENDERGAST:2014 . With the development of monitoring approaches, the sensors can facilitate data collection in real-time concerning bed elevation and other flow-related information. Compared to traditional databases sheppard2011scour , the sensor-collected data are richer and more granular. Therefore, the monitoring data can also be utilized for scour depth prediction and providing early-warning information.
Yousefpour et al. phase1 ; phase2 pioneered the idea of incorporating real-time scour monitoring data to develop real-time scour forecast models using LSTMs. They incorporated the monitoring data in Alaska (2006 to 2017) collected by the U.S. Geological Survey (USGS) in partnership with the Departments of Transport (DOTs) in the US. Specifically, historical time series of flow depth and bed elevation were used to train the LSTM models to predict future bed elevation and upcoming scour depth. The proposed model showed promising results in forecasting bed-elevation changes for at least seven days in advance. Compared to traditional empirical equations, these time-series forecasting models showed to yield more reliable predictions of upcoming scour trends and maximum scour depth.
Chang et al. chang2014pier presented a real-time monitoring system, utilizing micro-cameras mounted on bridge piers, in conjunction with image processing and pattern recognition methodologies, to monitor variations in bed levels. This approach was validated through rigorous laboratory experiments, illustrating that the image data can be transformed to capture the progressive alterations in scour depth. Typically, monitoring sensors are installed on bridges deemed high risk.
@@ 4.2 Feature Engineering
We also engineered another feature, the equivalent velocity $X_{\textit{eVelocity}}$, which is directly proportional to the flow velocity following Equations. 1 to  4. By assuming a constant channel equivalent cross-sectional width $\alpha$, the channel cross section area can be estimated by $\alpha$ x $d$, where $d$ is the flow depth at a bridge pier equal to the difference between Sonar and Stage elevations (Equation  1).
@@
$X_{\textit{discharge}}$ $=d*\alpha*velocity$ (2)
$X_{\textit{eVelocity}}$ $=\alpha*velocity$ (3)
@@
year-sin $=\sin(\textit{TS}*(2\pi/\textit{year}))$ (5)
year-cos $=\cos(\textit{TS}*(2\pi/\textit{year}))$ (6)
@@ 4.3.1 LSTM Models
Long Short Term Memory (LSTM)  hochreiter1997long is a specialised recurrent neural network (RNN) that is capable of modelling very large input and output sequences and does not suffer from the gradient vanishing problem - a typical problem in RNNs  LSTM1 ; LSTM2 . The principle analogy of LSTM is to remember the relevant short and long-standing information and discard the irrelevant information. Three different gates, i.e., Forget gate ($\Gamma^{f}$), Input gate ($\Gamma^{i}$) and Output gate ($\Gamma^{o}$), enable the LSTM model to achieve this goal as shown in Fig.  1 (a). In these gates, the degree of relevance of information is quantified using a sigmoid ($\sigma$) function with a value between 00 and $1$. The larger the value, the higher the relevance. Equations (7 - 12) show how the hidden unit output ($a_{t}$) and LSTM memory cell output ($c_{t}$) are recurrently computed using these gates. $W$ and $b$ represent the weight and bias matrices.
@@
$c_{t}$ $=\Gamma^{i}*\tilde{c_{t}}+\Gamma^{f}*c_{t-1}$ (11)
$a_{t}$ $=\Gamma^{o}*\tanh(c_{t})$ (12)
[Uncaptioned image] [Uncaptioned image]
(b) (c)
Table 1: Figure (a) shows the underlying configurations of (i) Forget, (ii) Update, and (iii) Output gates in an LSTM memory unit, (b) Single-Shot and (c) Feedback variants
@@
The following three variants introduced in phase2 are implemented in this study:
@@
    Single-Shot (ss): In this model, one LSTM layer has been followed by one Dense and one Reshape layer [see Fig.  1 (b)]. This model outputs the predictions for all the timesteps over the forecast window ($h$) at once. Fig.  4 shows the architecture of the Single-Shot LSTM model for the multivariate time series forecast.
@@
    Two-Layer, Single-Shot (ss2): Two LSTM layers are stacked together to add more complexity to the LSTM network so that it captures more variability of the input features. The rest of the setting remains the same as Single-Shot.
@@
    Feedback (Auto-Regressive: fb): In this variation, we hypothesize that the model’s prediction accuracy would get improved, if we input the predicted value (as opposed to the actual value) of each timestep within the label/forecast window for the prediction of the next timestep [^2] Google Tensorflow: https://www.tensorflow.org/tutorials/structured_data/time_series. Therefore, in this variant, LSTM predicts one timestep ahead across the entire forecast window. Fig.  1 (c) illustrates the feedback process. Only one LSTM layer is used, similar to the Single-Shot model.
@@ 4.3.2 CNN Models
CNN models proposed in  earlyCNN are widely used in computer vision for tasks like image classification  CNN_AlexNet and object detection  CNNobjectdetection . They offer excellent performance with lower computational requirements. CNNs also excel in time series modelling tasks  CNN_timeseries ; CNNintimeseries ; TCN . Here, we investigated three CNN variants for multi-variate time-series scour forecasting (see Fig.  5).
@@
In time-series forecast, a pivotal element of the CNN model is the convolution process—a filter sliding over input time-sequence for feature extraction as depicted in Fig.  6. Equation  13 defines the standard convolution operation ($C_{\text{vanilla}}$) as a dot product between a filter ($F$) and a time-sequence ($S$) of lengths $k$ and $l$ respectively. Because of the weight-sharing property of a CNN model, the entire time sequence is repeatedly convolved by the same filter. Fig.  6 (a) illustrates this vanilla convolution using a $3$-step filter on a $9$-step sequence with a stride of $1$. Convolution results in a reduced-length sequence of $7$ steps ($l-k+1=9-3+1=7$), which leads to information loss. To prevent this loss, padded convolution ($C_{\text{padded}}$) operates on an extended sequence $S^{\prime}$, created by padding zeros to the edges of $S$ [see Equation  14 and Fig.  6 (b)]. This ensures the input and convolved time sequences have the same length.
$C_{vanilla}$ $:=F*S_{t:t+k}\;|\;\forall t\in[1,l-k]$ (13)
$C_{padded}$ $:=F*S^{\prime}_{t:t+k}\;|\;\forall t\in[0,l-k+1]$ (14)
$C_{casual}$ $:=F*S^{\prime}_{t-k+1:t}\;|\;\forall t\in[1,l]$ (15)
$C_{dilated-casual}$ $:=F^{\prime}*S^{\prime}_{t-(k+d-1)+1:t}\;|\;\forall t\in[1,l]$ (16)
[Uncaptioned image] [Uncaptioned image]
@@
[Uncaptioned image] [Uncaptioned image]
@@
Figure 6: Illustration of different types of CNN convolutions for a $9$-length time-series. Figure (a) shows a standard (vanilla) convolution process by a $3$-length filter. Figure (b) shows the padded convolution. Figures (c) and (d) show the causal and dilated-causal convolutions by a $2$-length filter, respectively. The dilation factor is considered to be $1$ and $2$ for the bottom and top layers in Figure (d).
Casual convolution ($C_{casual}$) is often a preferable algorithm in time-series forecasting to avoid information leakage from the future  CNN_casualdilated . It precedes the input sequence with $k$-1 zeroes before convolution, as illustrated in Fig.  6 (c). Casual convolution would require a large filter or a deep network when it needs to examine a large number of past observations. Dilated Convolution overcomes this drawback by introducing a dilation factor $d$ into the filter so that convolution can be applied on an input area larger than the original filter size, i.e., capturing larger contextual information. The modified filter $F^{{}^{\prime}}$ is obtained by dilating the original filter with ($d-1$) zeros as shown in Fig.  6 (d). Dilation is performed using $d=2$. The positions of zero correspond to the skipped timesteps in the input sequence. Both the casual (in Equation  15) and dilated casual convolutions ($C_{dilated-casual}$) in Equation  16 output a sequence with the same length of the input. In the following, we introduce three CNN model configurations to perform the multivariate scour forecast:
@@ 5.1 Evaluation of Random Policies for HP Tuning (Experiment 1)
Fig.  7 demonstrates the results of HP tunning using grid-search and random-search policies for the $18$ scour forecast models. Top-3 models with the smallest average MAE from grid-search (GS) served as benchmarks when evaluating random-search (RS) policies. Since the distribution of MAE values is not heavily skewed as observed in the boxplots, the mean values are used to rank the models in Fig.  7 (a). It is noted that higher ranked models (top-$5$) would remain the same with median criteria. Note that all the reported MAE values throughout the paper are in m (of elevation).
@@
The meanMAE and medianMAE random-search policies show similar trends in lower sample sizes (35% and 50%), but as expected, they deviate by increasing the sample size. With a growing sample size, the mean-MAE estimations seem to converge towards grid-search estimates. The bagging policy also fails to find the top-ranked models when using small sampling sizes; however, with increasing sample size (67%), this method shows to be comparatively more robust and less sensitive to the poor-performing models compared to the meanMAE and medianMAE policies (e.g. models 12 to 18 show low ranking based on bagging but rank high in mean and median policies). Furthermore, the meanMAE and bagging policies seem to agree with 2 out of 3 top selected configurations, whereas the medianMAE policy fails to identify any of the top-performing models [Fig.  7 (e)].

### v2

# Application of Long-Short Term Memory and Convolutional Neural Networks for Real-Time Bridge Scour Prediction

@@ Abstract
Scour around bridge piers is a critical challenge for infrastructures around the world. In the absence of analytical models and due to the complexity of the scour process, it is difficult for current empirical methods to achieve accurate predictions. In this paper, we exploit the power of deep learning algorithms to forecast the scour depth variations around bridge piers based on historical sensor monitoring data, including riverbed elevation, flow elevation, and flow velocity. We investigated the performance of Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN) models for real-time scour forecasting using data collected from bridges in Alaska and Oregon from 2006 to 2021. The LSTM models achieved mean absolute error (MAE) ranging from 0.1m to 0.5m for predicting bed level variations a week in advance, showing a reasonable performance. The Fully Convolutional Network (FCN) variant of CNN outperformed other CNN configurations, showing a comparable performance to LSTMs with significantly lower computational costs. We explored various innovative random-search heuristics for hyperparameter tuning and model optimisation which resulted in reduced computational cost compared to grid-search method. The impact of different combinations of sensor features on scour prediction showed the significance of the historical time series of scour for predicting upcoming events. Overall, this study provides a greater understanding of the potential of Deep Learning algorithms for real-time scour prediction and early warning for bridges with distinct geology, geomorphology and flow characteristics.
@@
Scour Prediction, Deep Learning, Time Series Forecasting, Hyperparameter Tuning, Long-short Term Memory Network, Convolutional Neural Network
@@ 1 Introduction
Artificial Intelligence (AI) and Machine Learning (ML) for scour depth estimation have been investigated by a range of algorithmic strategies, such as evolutionary computing, fuzzy logic, artificial neural networks, support vector machines, and decision trees (see section  2). These proposed techniques have shown promising results in estimating maximum scour depth, often outperforming traditional empirical equations. However, the performance of these models is still limited by the availability and quality of training data, which are often scarce and may not cover a wide range of geological, hydraulic and geomorphological conditions. Furthermore, these models are typically trained to predict a maximum scour depth for a given flow discharge and cannot be used for dynamic scour prediction and real-time forecasting.
Yousefpour et al. phase1 ; phase2 pioneered the application of deep learning (DL) solutions, using Long Short Term Memory (LSTM) networks hochreiter1997long for real-time scour forecasting. They leveraged the LSTM’s superiority in temporal pattern recognition and capturing the underlying physics without direct feature extraction. They used historical scour monitoring data in Alaska to train DL models and predict future bed elevation variations and upcoming scour events. Fig. 1 shows the process of local scour around a bridge pier; scour happens when sediments are eroded and washed away by strong vortices in interaction with the bridge substructure, often accelerating during floods. Stage (flow/water level) and sonar (bed elevation) sensors are commonly used sensors installed on the bridge piers to monitor scour by collecting real-time flow and bed elevation data. The variants of LSTM models developed in phase2 showed promising performance in forecasting the scour depth a week in advance, for case-study bridges in Alaska. However, there remain research questions about the capability of these algorithms to be upscaled for implementation in practice and deployment in different locations with variable meteorological, geological, geomorphological, and flow conditions.
@@
In this study, we addressed fundamental research questions about the efficiency and generalization of the LSTM scour forecast models in diverse spatio-temporal domains. The LSTM models are compared against temporal CNN as an alternative deep learning solution, in terms of performance and computational cost. Random-Search (RS) heuristics for hyperparameter (HP) tuning are evaluated  RS1 to explore alternate computationally efficient methods to Grid-Search (GS) used in  phase2 for model optimisation. We also investigated the impacts of monitored features (input parameters), including stage, sonar, and velocity (discharge), on the performance of scour forecast models.
To summarise, four research questions were addressed in this study:
@@
(RQ2) How do other competitor models, including CNN with vanilla and dilated convolutions, compare to LSTMs for the prediction of scour?
@@
We carefully designed experiments and developed heuristics to investigate the aforementioned research questions as discussed in the following sections. The DL models were trained on the University of Melbourne’s Spartan high-performance computing (GPU and CPU) cluster. The models were developed in Python using Tensorflow library tensorflow2015-whitepaper .
@@ 2.2 Real-time Scour Monitoring and Forecast
Although the worst-case scenario (usually a 1-in-100-year flood) is considered for estimating maximum scour depth in bridge design, many uncertainty factors could still lead to scour-related bridge failures: Larger flood events can occur due to extreme weather conditions and climate change; complexities in the scour process, particularly river sediment material interaction with flow which can result in larger scour depth; older bridges could fall behind the best recommended practices for scour estimation khandel2019integrated ; lamb2017vulnerability . Considering all these factors, the transport authorities in the past two decades have been resorting to regular monitoring to manage scour, including regular inspections and surveys as well as advanced remote sensing.
In recent years, real-time scour monitoring sensors have become a prominent solution installed on many large-scale and critical bridges lagasse1997instrumentation ; briaud2011realtime ; lebbe2014failure ; PRENDERGAST:2014 . With the development of monitoring approaches, the sensors can facilitate data collection in real-time concerning bed elevation and other flow-related information. Compared to traditional databases sheppard2011scour , the sensor data collected is richer and more granular. Therefore, the monitoring data can also be utilized for scour depth prediction and providing early-warning information.
Yousefpour et al. phase1 ; phase2 pioneered the idea of incorporating real-time scour monitoring data to develop real-time scour forecast models using LSTMs. They incorporated monitoring data in Alaska (2006 to 2017) collected by the US Geological Survey (USGS) in partnership with the Departments of Transport (DOTs) in the United States. Specifically, historical time series of flow depth and bed elevation were used to train the LSTM models to predict future bed elevation and upcoming scour depth. The proposed model showed promising results in forecasting bed-elevation changes for at least seven days in advance. Compared to traditional empirical equations, these time series forecasting models yielded more reliable predictions of the upcoming scour trends and maximum depth of scour.
Chang et al. chang2014pier presented a real-time monitoring system, using micro-cameras mounted on bridge piers, in conjunction with image processing and pattern recognition methodologies, to monitor variations in bed levels. This approach was validated through rigorous laboratory experiments, illustrating that the image data can be transformed to capture the progressive alterations in scour depth. Typically, monitoring sensors are installed on bridges deemed high risk.
@@ 4.2 Feature Engineering
We also engineered another feature, the equivalent velocity $X_{\textit{eVelocity}}$, which is directly proportional to the flow velocity ($V$) following Equations. 1 to  4. By assuming a constant channel equivalent cross-sectional width $\alpha$, the channel cross section area can be estimated by $\alpha$ x $d$, where $d$ is the flow depth at a bridge pier equal to the difference between Sonar and Stage elevations (Equation  1).
@@
$X_{\textit{discharge}}$ $=\alpha dV$ (2)
$X_{\textit{eVelocity}}$ $=\alpha V$ (3)
@@
year-sin $=\sin(\textit{TS}.2\pi/\textit{year})$ (5)
year-cos $=\cos(\textit{TS}.2\pi/\textit{year})$ (6)
@@ 4.3.1 LSTM Models
Long Short Term Memory hochreiter1997long is a specialised recurrent neural network (RNN) that is capable of modelling very large input and output sequences and does not suffer from the gradient vanishing problem - a typical problem in RNNs LSTM1 ; LSTM2 . The principle analogy of LSTM is to remember the relevant short and long-standing information and discard the irrelevant information. Three different gates, i.e., Forget gate ($\Gamma^{f}$), Input gate ($\Gamma^{i}$) and Output gate ($\Gamma^{o}$), enable the LSTM model to achieve this goal as shown in Fig.  4 (a). In these gates, the degree of relevance of information is quantified using a sigmoid ($\sigma$) function with a value between 00 and $1$. The larger the value, the higher the relevance. Equations (7 - 12) show how the hidden unit output ($a_{t}$) and LSTM memory cell output ($c_{t}$) are recurrently computed using these gates. $W$ and $b$ represent the weight and bias matrices.
@@
$c_{t}$ $=\Gamma^{i}\tilde{c_{t}}+\Gamma^{f}c_{t-1}$ (11)
$a_{t}$ $=\Gamma^{o}\tanh(c_{t})$ (12)
Figure 4: (a) LSTM memory unit, (b) Single-Shot and (c) Feedback variants
@@
Fig.  5 shows the architecture of the LSTM models for the multivariate time series forecasting of scour. The following three variants introduced in phase2 are also implemented in this study:
@@
    Single-Shot (ss): In this model, one LSTM layer has been followed by one Dense and one Reshape layer. This model outputs the predictions for all the timesteps over the forecast window at once, as shown in Fig. 4(b).
@@
    Two-Layer, Single-Shot (ss2): Two LSTM layers are stacked together in this variant. The rest of the setting remains the same as Single-Shot.
@@
    Feedback (fb): In this variation, we hypothesize that the model’s prediction accuracy would get improved, if we input the predicted value (as opposed to the actual value) of each timestep for the prediction of the next timestep. Therefore, in this variant, LSTM predicts one timestep ahead across the entire forecast window. Fig. 4 (c) illustrates the feedback process. The feedback variant also has one LSTM layer, similar to the Single-Shot model.
@@ 4.3.2 CNN Models
CNN models proposed in  earlyCNN are widely used in computer vision for tasks like image classification  CNN_AlexNet and object detection  CNNobjectdetection . They offer excellent performance with lower computational requirements. CNNs also excel in time series modelling tasks  CNN_timeseries ; CNNintimeseries ; TCN . We investigated three CNN variants for multi-variate time-series scour forecasting with an architecture as shown in Fig. 6).
@@
In time-series forecast, a pivotal element of the CNN model is the convolution process—a filter sliding over input time-sequence for feature extraction as depicted in Fig.  7. Equation  13 defines the standard convolution operation ($C_{\text{vanilla}}$) as a dot product between a filter ($F$) and a time-sequence ($S$) of lengths $k$ and $l$ respectively. Because of the weight-sharing property of a CNN model, the entire time sequence is repeatedly convolved by the same filter. Fig.  7 (a) illustrates this vanilla convolution using a $3$-step filter on a $9$-step sequence with a stride of $1$. Convolution results in a reduced-length sequence of $7$ steps ($l-k+1=9-3+1=7$), which leads to information loss. To prevent this loss, padded convolution ($C_{\text{padded}}$) operates on an extended sequence $S^{\prime}$, created by padding zeros to the edges of $S$ [see Equation  14 and Fig.  7 (b)]. This ensures the input and convolved time sequences have the same length.
$C_{vanilla}$ $:=F.S_{t:t+k}\;|\;\forall t\in[1,l-k]$ (13)
$C_{padded}$ $:=F.S^{\prime}_{t:t+k}\;|\;\forall t\in[0,l-k+1]$ (14)
$C_{casual}$ $:=F.S^{\prime}_{t-k+1:t}\;|\;\forall t\in[1,l]$ (15)
$C_{dilated-casual}$ $:=F^{\prime}.S^{\prime}_{t-(k+d-1)+1:t}\;|\;\forall t\in[1,l]$ (16)
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
@@
Figure 7: Illustration of different types of CNN convolutions for a $9$-length time-series: (a) standard (vanilla) convolution process by a $3$-length filter, (b) padded convolution, (c) casual convolution and (d) dilated-causal convolution by a $2$-length filter.
Casual convolution ($C_{casual}$) is often a preferable algorithm in time-series forecasting to avoid information leakage from the future  CNN_casualdilated . It precedes the input sequence with $k$-1 zeroes before convolution, as illustrated in Fig. 7 (c). Casual convolution would require a large filter or a deep network when it needs to examine a large number of past observations. Dilated Convolution overcomes this drawback by introducing a dilation factor $d$ into the filter so that convolution can be applied on an input area larger than the original filter size, i.e., capturing larger contextual information. The modified filter $F^{{}^{\prime}}$ is obtained by dilating the original filter with ($d-1$) zeros as shown in Fig.  7 (d). Dilation is performed using $d=2$. The positions of zero correspond to the skipped timesteps in the input sequence. Both the casual (in Equation  15) and dilated casual convolutions ($C_{dilated-casual}$) in Equation  16 output a sequence with the same length of the input. In the following, we introduce three CNN model configurations to perform the multivariate scour forecast:
@@ 5.1 Evaluation of Random Policies for HP Tuning (Experiment 1)
Fig.  8 demonstrates the results of HP tunning using grid-search and random-search policies for the $18$ scour forecast models. Top-3 models with the smallest average MAE from grid-search (GS) served as benchmarks when evaluating random-search (RS) policies. Since the distribution of MAE values is not heavily skewed as observed in the boxplots, the mean values are used to rank the models in Fig.  8 (a). It is noted that higher ranked models (top-$5$) would remain the same with median criteria. Note that all the reported MAE values throughout the paper are in m (of elevation).
@@
The meanMAE and medianMAE random-search policies show similar trends in lower sample sizes (35% and 50%), but as expected, they deviate by increasing the sample size. With a growing sample size, the mean-MAE estimations seem to converge towards grid-search estimates. The bagging policy also fails to find the top-ranked models when using small sampling sizes; however, with increasing sample size (67%), this method shows to be comparatively more robust and less sensitive to the poor-performing models compared to the meanMAE and medianMAE policies (e.g. models 12 to 18 show low ranking based on bagging but rank high in mean and median policies). Furthermore, the meanMAE and bagging policies seem to agree with 2 out of 3 top selected configurations, whereas the medianMAE policy fails to identify any of the top-performing models [Fig.  8 (e)].